Exercise 14-09 (Part Level Submission)

On June 30, 2020, Larkspur Company issued $3,660,000 face value of 13%, 20-year bonds at $3,935,340, a yield of 12%. Larkspur uses the effective-interest method to amortize bond premium or discount. The bonds pay semiannual interest on June 30 and December 31.

Collapse question part

(a)

Prepare the journal entries to record the following transactions. (Round answer to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)
(1) The issuance of the bonds on June 30, 2020.
(2) The payment of interest and the amortization of the premium on December 31, 2020.
(3) The payment of interest and the amortization of the premium on June 30, 2021.
(4) The payment of interest and the amortization of the premium on December 31, 2021.

No.

Date

Account Titles and Explanation

Debit

Credit

(1)

June 30, 2020

(2)

December 31, 2020

(3)

June 30, 2021

(4)

December 31, 2021

SHOW LIST OF ACCOUNTS

LINK TO TEXT

Attempts: 0 of 3 used

SAVE FOR LATER

SUBMIT ANSWER

Expand question part

(b)

The parts of this question must be completed in order. This part will be available when you complete the part above.

Expand question part

(c)

The parts of this question must be completed in order. This part will be available when you complete the part above.
Answer #1
Date Account Titles and Explanation Debit Credit
a June 30, 2020 Cash 3935340
      Bonds payable 3660000
      Premium on Bonds payable 275340
b December 31, 2020 Interest expense 236120 =3935340*12%/2
Premium on Bonds payable 1780
     Cash 237900 =3660000*13%/2
c June 30, 2021 Interest expense 236014 =(3935340-1780)*12%/2
Premium on Bonds payable 1886
     Cash 237900 =3660000*13%/2
d December 31, 2021 Interest expense 235900 =(3935340-1780-1886)*12%/2
Premium on Bonds payable 2000
     Cash 237900 =3660000*13%/2